OpenStudy (anonymous):

An 8.10 kilogram box of mass is kept at a height of 0.99 meters above ground level. What is the potential energy of the box? (given g = 9.8 meters/second2) A. 68 joules B. 79 joules C. 81 joules D. 99 joules E. 1.0 × 102 joules

OpenStudy (ghazi):

Potential energy possessed by the box is \[P.E= m*g*H=8.10*9.8*0.99=79 J\]